OpenStudy (anonymous):

You are a manufacturer of a certain type of watch, and you’ve calculated that it costs $600 to make 35 watches, and it costs $715 to make 50 watches. Assuming that the relationship between cost and the number of watches you make (volume) is linear, write the equation that relates cost to the number of watches you make and use the equation to calculate how much it would cost to produce 42 watches. Round your answer to the nearest dollars and cents. Be sure to use the dollar sign and the decimal point for the correct answer

OpenStudy (anonymous):

Linear relationship between cost C, & number of watches W so C(W) = aW + b C(35) = 600 = a*35 + b C(50) = 175 = a*50 + b solve the simultaneous equations, then find C(42). if you need more help than that, say so.